### v3

@@ 7 Discussion
Consider any two actions $a_{k},a_{l}$ and for any $\theta$, denote $A^{\theta}_{k,l}=u^{\theta}_{k}-u^{\theta}_{l}$. As noted above, the proof of sufficiency relies on the main result of Caplin and Martin, 2015, who show that rationalizability with a Bayesian expected utility DM is equivalent to the existence of a solution to a certain system of inequalities (see Appendix A. BEU-Rationalizability). We show that MLR property ensures that a solution exists, which is also single crossing.
@@
where one of two inequalities is strict. Suppose that $u_{k}^{\theta_{N}}-u_{l}^{\theta_{N}}<0$. Then, SCP requires that $u_{k}^{\theta}-u_{l}^{\theta}<0$ for all $\theta$ - however, this would imply that numbers $u^{\theta}_{k},u^{\theta}_{l}$ cannot be a solution to the inequality $\sum_{\theta}q^{\theta}_{k}[u^{\theta}_{k}-u^{\theta}_{l}]\geq 0$. The analyst must then reject the single crossing property. Similarly, if $u_{k}^{\theta_{1}}-u_{l}^{\theta_{1}}>0$, then $u_{k}^{\theta}-u_{l}^{\theta}>0$ for every $\theta$ and therefore the numbers $u^{\theta}_{k},u^{\theta}_{l}$ cannot be a solution to the inequality $\sum_{\theta}q^{\theta}_{l}[u^{\theta}_{k}-u^{\theta}_{l}]\leq 0$. However, under uncertainty, almost all of the bite of single crossing property rests upon the information structure being MLR ordered. Indeed, one implication of the result of Athey, 2002 is that, if the information structure is not MLR ordered, then there exists a utility function satisfying the SCP such that MCS does not hold. Thus, under uncertainty, most of the bite of SCP comes when combined with an MLR ordered information structures.
Furthermore, without uncertainty, the only observable implication is monotonicity of chosen actions with respect to the exogenous covariate (e.g. known state), as shown in Lazzati, Quah and Shirai, 2018. As pointed out above, the single crossing property is an ordinal generalization of the property of increasing differences. The latter is intimately related to supermodularity. [^24] A function $f:X\times\Theta\rightarrow\mathbb{R}$ is said to be supermodular if $f(x,\theta)+f(x^{\prime},\theta^{\prime})\leq f((x,\theta)\vee(x^{\prime},\theta^{\prime}))+f((x,\theta)\wedge(x^{\prime},\theta^{\prime}))$, where $X\times\Theta$ is ordered according to the product ordering induced by the orderings over the sets $X$ and $\Theta$. If the functions $x\mapsto f(x,\theta)$ and $\theta\mapsto f(x,\theta)$ are both supermodular for any $\theta$ and $x$ respectively, and if $f$ has increasing differences then $f$ is supermodular. See section 2.6.1 of Topkis, 1998. Interestingly, Chambers and Echenique, 2009 argue that quasi-supermodularity, the ordinal generalization of the notion of supermodularity is indistinguishable from the latter if the underlying domain is finite and the function is assumed to be monotonic. In several settings, therefore, this property is not refutable. A novelty of the results reported here is thus establishing an observational non-equivalence between monotonicity and single crossing property for certain kinds of data.
@@
Rambachan, 2024 shows how to combine the NIAS inequalities with known methods of partial identification to obtain a novel theory of identifying prediction mistakes with stochastic choice data. Our results can thus be possibly used to further sharpen the identification bounds by imposing some shape restrictions.
@@ Appendix A. BEU-Rationalizability
This section briefly presents the result of Caplin and Martin, 2015 since the latter is the result on which our paper builds. A BEU-rationalization of a dataset is $\mathcal{D}$ is a tuple $\langle u,\pi,C\rangle$ such that conditions 242-4 in Definition 4 hold.
Theorem 5 (Caplin and Martin, 2015, Theorem 1).

### v4

@@ 7 Discussion
Consider any two actions $a_{k},a_{l}$ and for any $\theta$, denote $A^{\theta}_{k,l}=u^{\theta}_{k}-u^{\theta}_{l}$. As noted above, the proof of sufficiency relies on the main result of caplin2015testable, who show that rationalizability with a Bayesian expected utility DM is equivalent to the existence of a solution to a certain system of inequalities (see Appendix A. BEU-Rationalizability). We show that MLR property ensures that a solution exists, which is also single crossing.
@@
where one of two inequalities is strict. Suppose that $u_{k}^{\theta_{N}}-u_{l}^{\theta_{N}}<0$. Then, SCP requires that $u_{k}^{\theta}-u_{l}^{\theta}<0$ for all $\theta$ - however, this would imply that numbers $u^{\theta}_{k},u^{\theta}_{l}$ cannot be a solution to the inequality $\sum_{\theta}q^{\theta}_{k}[u^{\theta}_{k}-u^{\theta}_{l}]\geq 0$. The analyst must then reject the single crossing property. Similarly, if $u_{k}^{\theta_{1}}-u_{l}^{\theta_{1}}>0$, then $u_{k}^{\theta}-u_{l}^{\theta}>0$ for every $\theta$ and therefore the numbers $u^{\theta}_{k},u^{\theta}_{l}$ cannot be a solution to the inequality $\sum_{\theta}q^{\theta}_{l}[u^{\theta}_{k}-u^{\theta}_{l}]\leq 0$. However, under uncertainty, almost all of the bite of single crossing property rests upon the information structure being MLR ordered. Indeed, one implication of the result of athey2002monotone is that, if the information structure is not MLR ordered, then there exists a utility function satisfying the SCP such that MCS does not hold. Thus, under uncertainty, most of the bite of SCP comes when combined with an MLR ordered information structures.
Furthermore, without uncertainty, the only observable implication is monotonicity of chosen actions with respect to the exogenous covariate (e.g. known state), as shown in lazzati2018nonparametric. As pointed out above, the single crossing property is an ordinal generalization of the property of increasing differences. The latter is intimately related to supermodularity. [^24] A function $f:X\times\Theta\rightarrow\mathbb{R}$ is said to be supermodular if $f(x,\theta)+f(x^{\prime},\theta^{\prime})\leq f((x,\theta)\vee(x^{\prime},\theta^{\prime}))+f((x,\theta)\wedge(x^{\prime},\theta^{\prime}))$, where $X\times\Theta$ is ordered according to the product ordering induced by the orderings over the sets $X$ and $\Theta$. If the functions $x\mapsto f(x,\theta)$ and $\theta\mapsto f(x,\theta)$ are both supermodular for any $\theta$ and $x$ respectively, and if $f$ has increasing differences then $f$ is supermodular. See section 2.6.1 of topkis1998supermodularity. Interestingly, chambers2009supermodularity argue that quasi-supermodularity, the ordinal generalization of the notion of supermodularity is indistinguishable from the latter if the underlying domain is finite and the function is assumed to be monotonic. In several settings, therefore, this property is not refutable. A novelty of the results reported here is thus establishing an observational non-equivalence between monotonicity and single crossing property for certain kinds of data.
@@
rambachan2024identifying shows how to combine the NIAS inequalities with known methods of partial identification to obtain a novel theory of identifying prediction mistakes with stochastic choice data. Our results can thus be possibly used to further sharpen the identification bounds by imposing some shape restrictions.
@@ Appendix A. BEU-Rationalizability
This section briefly presents the result of caplin2015testable since the latter is the result on which our paper builds. A BEU-rationalization of a dataset is $\mathcal{D}$ is a tuple $\langle u,\pi,C\rangle$ such that conditions 242-4 in Definition 4 hold.
Theorem 5 (caplin2015testable, Theorem 1).